eclipsedstar (eclipsedstar):

A computer can perform c calculations in s seconds. How many minutes will it take the computer to perform k calculations? A. 60ks/c B. ks/c C. ks/60c D. 60c/ks

eclipsedstar (eclipsedstar):

@King.Void.

OpenStudy (king.void.):

Your answer will be found in the link below! https://answers.yahoo.com/question/index?qid=20130724024528AAeTc2l Good luck! Void!~

eclipsedstar (eclipsedstar):

o.O

eclipsedstar (eclipsedstar):

lol

eclipsedstar (eclipsedstar):

it's C

OpenStudy (king.void.):

Once again. ~sucks at math~

OpenStudy (anonymous):

Wow

eclipsedstar (eclipsedstar):

@wio I'm sorry...why "wow"? X)

OpenStudy (anonymous):

I just wonder if you understand how to do it

eclipsedstar (eclipsedstar):

nope, just need to know the formula, I'm still in 10th grade XD

eclipsedstar (eclipsedstar):

The teacher just added the random question somehow, extra credit XD

OpenStudy (anonymous):

I think you would set up a proportion... I wonder what they taught you so far.

eclipsedstar (eclipsedstar):

Well, I'm taking an online class, so...

eclipsedstar (eclipsedstar):

I can't exactly explain what I'm learning... :)

OpenStudy (anonymous):

The proportion here is:\[ \frac{\text{time}}{\text{calculations}}=\frac xk = \frac sc \implies x = \frac {ks}c \]But \(x\) is in seconds, so you divide by \(60\) to make it minutes...\[ \frac{ks}{60c} \]

eclipsedstar (eclipsedstar):

ohh ok... I see :D

eclipsedstar (eclipsedstar):

Thanks! :D

eclipsedstar (eclipsedstar):

Better than the link @King.Void. gave...hahaha jk